What is the graph 4y + 3x = -4
Found 2 solutions by ewatrrr, emargo19:
Answer by ewatrrr(10682)


Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
4y + 3x = -4 |solving for y to put into slope-intercept form
y = -(3/4)x - 1
y-intercept is (0,-1) slope of the line m = -(3/4) line slants to left m < 0
Note: slope tells us as y value changes by 3, x value changes by 4 in the opposite direction


Answer by emargo19(101)

Slope intercept equation (y=mx+b)
b = y-intercept
m=slope
4y + 3x = -4
4y = -3x - 4
y = -3/4x -1
b = -1 (Therefore y=-1)
Lets plot y=-1 and m=-3/4
I'll start by plotting y=-1
From this point, I go down -3 on the y-axis and over 4 to the left on x-axis: